Anna has 12 bills in her wallet, some $5 and some $10. The total value of the bills is $100. How many of each bill does Anna have?
Answer by stanbon(75887)

Anna has 12 bills in her wallet, some $5 and some $10. The total value of the bills is $100. How many of each bill does Anna have?
------
Let number of $5 bill be x ; their value is 5x dollars
Number of $10 bills is "12-x"/ value is 10(12-x)=120-10x dollars
-----------
EQUATION:
value + value = $100
5x+120-10x=100
-5x=-20
x=4 (# of $5 bills)
12-x=8 (# of $10 bills)
Cheers,
Stan H.